# Multi-Agent Quantum Reinforcement Learning using Evolutionary Optimization

### v3

@@ 7 CONCLUSION
Gradient-based training methods are, as of time of writing, not suitable for MAQRL due to problems with barren plateaus and vanishing gradients [Franz et al., 2022, Chen et al., 2022]. In this work, we presented a alternative approach to gradient based training methods for MAQRL. We build our approach upon the evolutionary optimization process used by [Chen et al., 2022], expanding it to Multi-Agent systems and different generational evolution strategies. We proposed three quantum approaches for MARL. All approaches use VQCs as replacement for neural networks. Two approach use recombination in addition to mutation and the third approach mutation only. For evaluation, we chose Coin Game as our testing environment due to it’s cooperative setting and relatively small observation space. As baselines, we used random agents and classical agents with neural networks, which were also trained using the evolutionary algorithm. To achieve a fair comparison, we chose a neural network with a similar amount of parameters, as well as one with a hidden layer size of $64\times 64$.

### v4

@@ 7 CONCLUSION
Gradient-based training methods are, as of time of writing, not suitable for MAQRL due to problems with barren plateaus and vanishing gradients [Franz et al., 2022, Chen et al., 2022].In this work, we presented an alternative approach to gradient based training methods for MAQRL. For our gradient-free approach used in the experiments, barren plateaus did not arise; however, it cannot be ruled out that such issues may occur in gradient-free methods under different conditions or in more complex scenarios [Larocca et al., 2024]. We build our approach upon the evolutionary optimization process used by [Chen et al., 2022], expanding it to Multi-Agent systems and different generational evolution strategies. We proposed three quantum approaches for MARL. All approaches use VQCs as replacement for neural networks. Two approach use recombination in addition to mutation and the third approach mutation only. For evaluation, we chose Coin Game as our testing environment due to it’s cooperative setting and relatively small observation space. As baselines, we used random agents and classical agents with neural networks, which were also trained using the evolutionary algorithm. To achieve a fair comparison, we chose a neural network with a similar amount of parameters, as well as one with a hidden layer size of $64\times 64$.